# How Does the Smoothness Approximation Method Facilitate Generalization for Federated Adversarial Learning?

### v1

@@ Conclusion
Appendix A Experiments Setups And Additional Experiments
Appendix B VFAL Algorithms
Algorithm 1 VFAL of SSA.
@@ Appendix B VFAL Algorithms
3:     Sample $\mathbf{u}^{t}\in\mathbb{R}^{d}$ uniformly from a unit sphere in $\mathbb{R}^{d}$ and send it to client $i$.
4:     for $k=0,1,...,K-1$ do
@@
Algorithm 3 VFAL of OPSA.
@@
Appendix C Useful Lemmas
@@
$\|\nabla R_{i}(\theta_{t})-\nabla R(\theta_{t})\|\leq(\xi+6L)D_{i}$
@@
Under Assumptions 1, we have
@@
$h(\theta_{1})-h(\theta_{2})\leq\nabla h(\theta^{\prime})^{T}(\theta_{1}-\theta_{2})+\frac{\beta}{2}\|\theta_{1}-\theta_{2}\|^{2}+\xi\|\theta_{1}-\theta_{2}\|$
$\|G_{\alpha,z}(\theta_{1})-G_{\alpha,z}(\theta_{2})\|\leq(1+\eta\beta)\|\theta-\theta^{\prime}\|+\eta\xi$
@@
Appendix D Generalization Analysis of VFAL
@@
There exists $\sigma>0$ such that for any $\theta$ and $i\in[m],\mathbb{E}_{z_{i}}[\nabla h(\theta,z_{i})]=\nabla R_{i}(\theta)$ and $\mathbb{E}\|\nabla h(\theta,z)-\nabla R_{i}(\theta)\|\leq\sigma^{2}$
@@
For simplicity, we use $g(\theta)$ to denote stochastic gradient of loss function.
@@
$=$ $-\frac{K\eta_{t}}{2}\mathbb{E}\|\nabla R(\theta_{t})\|^{2}-\frac{\eta_{t}}{2m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\nabla R_{i}(\theta^{t+1}_{i,k})\|^{2}+\frac{\eta_{t}\beta^{2}}{m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\theta^{t+1}_{i,k}-\theta_{t}\|^{2}+K\xi^{2}\eta_{t}$
@@
$\leq$ $\frac{\beta K\eta_{t}^{2}\sigma^{2}}{m}+\dfrac{\beta K\eta_{t}^{2}}{m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\nabla R_{i}(\theta^{t+1}_{i,k})\|^{2}$
@@
$+\frac{\beta K\eta_{t}^{2}\sigma^{2}}{m}-\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\dfrac{\eta_{t}}{2m}\left(1-2\beta K\eta_{t}\right)\mathbb{E}\|\nabla R_{i}(\theta_{i,k}^{t+1})\|^{2}$
@@
$\frac{1}{m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\theta^{t+1}_{i,k}-\theta_{t}\|^{2}\leq\frac{9}{2}K^{2}\eta_{t}^{2}\sigma^{2}+6K^{3}\eta_{t}^{2}\xi^{2}+\sum\limits_{i=1}^{m}\frac{3}{m}K^{3}\eta_{t}^{2}(\xi+6L)D_{i}^{2}+3K^{3}\eta_{t}^{2}\mathbb{E}\|\nabla R(\theta_{t})\|^{2}$ (12)
@@
$\frac{1}{m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\theta^{t+1}_{i,k}-\theta_{t}\|^{2}\leq\frac{9}{2}K^{2}\eta^{2}\sigma^{2}+6K^{3}\eta^{2}\xi^{2}+\sum\limits_{i=1}^{m}\frac{3}{m}K^{3}\eta^{2}(\xi+6L)^{2}D_{i}^{2}+3K^{3}\eta^{2}\mathbb{E}\|\nabla R(\theta_{t})\|^{2},$
@@
Suppose $h(\theta)$ is a $\xi$-approximately $\beta$-gradient Lipschitz and non-convex function w.r.t $\theta$.
$\mathbb{E}\|\theta_{i,k}-\theta_{t}\|\leq(1+\beta\eta_{t})^{K-1}K\eta_{t}(\xi+\sigma+(\xi+6L)D_{i}+\mathbb{E}\|\nabla R(\theta_{t})\|)$
@@
$\leq$ $(1+\beta\eta_{t})\mathbb{E}\|\theta_{i,k}-\theta_{t}\|+\eta_{t}(\xi+\sigma+(\xi+6L)D_{i}+\mathbb{E}\|\nabla R(\theta_{t})\|)$
@@
Suppose $h(\theta)$ is a $\xi$-approximately $\beta$-gradient Lipschitz and non-convex function w.r.t $\theta$, for $\eta_{t}\leq\frac{1}{K\beta(t+1)}$ , we have
@@
$+$ $e\sqrt{6\log T\Delta+6\log T(\xi^{2}+\xi L)+\dfrac{\log T\sigma^{2}\pi^{2}}{mK\beta}+\log T\zeta(3)c^{3}\beta^{2}(\frac{27\sigma^{2}}{K\beta}+\frac{36\xi^{2}}{\beta}+\frac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta})}$
@@
$\leq$ $(1+e)\left[\underbrace{\dfrac{K}{m}\sum\limits_{t=0}^{T-1}\sum\limits_{i=1}^{m}\eta_{t}(\sigma+\xi+(\xi+6L)D_{i})}_{\Lambda_{1}}+\underbrace{\dfrac{K}{m}\sum\limits_{t=0}^{T-1}\sum\limits_{i=1}^{m}\eta_{t}\mathbb{E}\|\nabla R(\theta_{t})\|}_{\Lambda_{2}}\right]$
@@
$\Lambda_{1}\leq$ $\dfrac{K}{m}\sum\limits_{t=0}^{T-1}\sum\limits_{i=1}^{m}\eta_{t}(\sigma+\xi+(\xi+6L)D_{i})\leq\log T(\sigma+\xi+(\xi+6L)D_{\max})$
@@
$+\dfrac{36\xi^{2}}{\beta(t+1)^{3}}+\dfrac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta(t+1)^{3}}$
@@
$=$ $\sqrt{6\log T\Delta+6\log T(\xi^{2}+\xi L)+\dfrac{\log^{2}T\sigma^{2}\pi^{2}}{mK\beta}+\log T\zeta(3)c^{3}\beta^{2}(\frac{27\sigma^{2}}{K\beta}+\frac{36\xi^{2}}{\beta}+\frac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta})}$
@@
$+$ $(1+e)\sqrt{6\log T\Delta+6\log T(\xi^{2}+\xi L)+\dfrac{\log T\sigma^{2}\pi^{2}}{mK\beta}+\log T\zeta(3)c^{3}\beta^{2}(\frac{27\sigma^{2}}{K\beta}+\frac{36\xi^{2}}{\beta}+\frac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta})}$
@@
$\leq$ $(1+\beta\eta_{t})\mathbb{E}\|\theta_{j,k}-\theta_{j,k}^{\prime}\|+\eta_{t}\xi$
@@
$=$ $e^{\beta K\eta_{t}}(\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+K\eta_{j,t}\xi)$
@@
$\leq$ $(1+\beta\eta_{t})\|\theta_{i,k}-\theta_{i,k}^{\prime}\|+\eta_{t}\xi$
@@
$\leq$ $(1+\beta\eta_{t})\|\theta_{i,k}-\theta_{i,k}^{\prime}\|+\eta_{t}\xi+\eta_{t}\|g_{i}(\theta_{i,k}^{\prime})-g_{i}^{\prime}(\theta_{i,k}^{\prime})\|$
@@
$\leq$ $(1+\beta\eta_{t})\mathbb{E}\|\theta_{i,k}-\theta_{i,k}^{\prime}\|+\eta_{t}\xi+\frac{2\eta_{t}}{n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|$
@@
$=$ $e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+K\eta_{t}e^{\beta K\eta_{t}}(\xi+\dfrac{2}{n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|)$
@@
$\leq$ $\dfrac{1}{m}\sum\limits_{j=1}^{m}e^{\beta K{\eta}_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+\dfrac{1}{m}\sum\limits_{j=1}^{m}e^{\beta K{\eta}_{t}}K\eta_{t}\xi+\dfrac{}{}+\dfrac{2K{\eta}_{t}e^{\beta K\eta_{t}}}{mn_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|$
@@
$\leq$ $e^{\frac{1}{(1+t)}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+e^{\frac{1}{(1+t)}}\sum\limits_{j=1}^{m}\frac{K\eta_{t}\xi}{m}+e^{\frac{1}{(1+t)}}\sum\limits_{i=1}^{m}\dfrac{2K\eta_{t}}{m^{2}n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|$
@@
$+$ $\frac{2\beta eT}{mn_{i}}\sqrt{6\log T\Delta+6\log T(\xi^{2}+\xi L)+\dfrac{\log T\sigma^{2}\pi^{2}}{mK\beta}+\log T\zeta(3)c^{3}\beta^{2}(\frac{27\sigma^{2}}{K\beta}+\frac{36\xi^{2}}{\beta}+\frac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta})},$
@@
$=$ $\mathcal{O}\left(\rho T\log T+\frac{T\sqrt{\log T\Delta}}{mn_{\mathrm{min}}}+\frac{T\log T(\rho+1)D_{\mathrm{max}}}{mn_{\mathrm{min}}}\right)$
@@
(Lin, Zheng, and Jordan (2022)Under assumption 1 and let $\ell_{\gamma}(\theta)=\mathbb{E}_{\mathbf{u}\sim\mathbb{P}}[\ell_{\rho}(\theta+\gamma u)]$ where $\mathbb{P}$ is an uniform distribution on an unit ball in $\mathcal{L}_{2}$ norm. We have
$\|\ell_{\gamma}(\theta)-\ell_{\rho}(\theta)\|\leq\gamma L$
$\|\ell_{\gamma}(\theta)-\ell_{\gamma}(\theta^{\prime})\|\leq L\|\theta-\theta^{\prime}\|$
$\|\nabla\ell_{\gamma}(\theta)-\nabla\ell_{\gamma}(\theta^{\prime})\|\leq\frac{cL\sqrt{d}}{\gamma}\|\theta-\theta^{\prime}\|$
@@
$=$ $L\|\theta-\theta^{\prime}\|$
@@
$\leq$ $\frac{L}{\mathrm{Vol}(\mathbb{B}_{\gamma}(\mathbf{0}))}\left(\mathrm{Vol}(\mathbb{B}_{\gamma}(\theta)\setminus\mathbb{B}_{\gamma}(\theta^{\prime}))+\mathrm{Vol}(\mathbb{B}_{\gamma}(\theta^{\prime})\setminus\mathbb{B}_{\gamma}(\theta))\right)$
@@
$=$ $\frac{2L}{c_{d}\gamma^{d}}\mathrm{Vol}(I)$
@@
$\mathrm{Vol}(I)=\mathrm{Vol}(\mathbb{B}_{\gamma}(\mathbf{0}))-2V_{cap}(\frac{1}{2}\|\theta-\theta^{\prime}\|)=c_{d}\gamma^{d}-2V_{cap}(\frac{1}{2}\|\theta-\theta^{\prime}\|).$
@@
$\leq$ $(1+\frac{cL\sqrt{d}}{\gamma}\eta)\|x-y\|$
@@
$\leq$ $(1+\eta)\frac{cL\sqrt{d}}{\gamma}\|\theta_{1}-\theta_{2}\|+2\eta v$
@@
If the Assumptions 1-2 hold and $\eta_{t}\leq\frac{\gamma}{6KcL\sqrt{d}}$, considering $\theta_{i,k+1}=\theta_{i,k}-\eta_{t}\nabla\ell_{\gamma,i}(\theta_{i,k})$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|R(\theta_{t})\|^{2}\leq\mathcal{O}(\frac{\sqrt{\sigma}}{\sqrt{mKT}}+\frac{((\xi+1)D_{\max})^{\frac{2}{3}}}{T^{\frac{2}{3}}}+\frac{\Delta}{T})$
@@
Theorem 3. Let $c\geq 0$ be a constant and the step size be chosen as $\eta_{t}\leq\frac{\gamma}{4K\sqrt{d}cL(t+1)}$. Under Assumption 1, the generalization bound $\varepsilon_{gen}$ with the randomized smoothness approximation satisfies:
@@
$\leq$ $(1+\frac{cL\sqrt{d}}{\gamma}\eta_{t})\mathbb{E}\|\theta_{j,k}-\theta_{j,k}^{\prime}\|+2\eta_{t}\mathbb{E}[v]$
@@
$=$ $e^{\frac{cL\sqrt{d}}{\gamma}K\eta_{t}}(\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+2K\eta_{t}\mathbb{E}[v])$
@@
$\leq$ $(1+\frac{cL\sqrt{d}}{\gamma}\eta_{t})\|\theta_{i,k}-\theta_{i,k}^{\prime}\|+2\eta_{t}\mathbb{E}[v]$

### v2

@@ Conclusion
Appendix A VFAL Algorithms
Algorithm 1 VFAL of SSA
@@ Appendix A VFAL Algorithms
3:     for $k=0,1,...,K-1$ do
4:        Sample $u_{i,k_{q}}\in\mathbb{P}^{d}$ uniformly from a unit sphere in $\mathbb{P}^{d}$.
@@
Algorithm 3 VFAL of OPSA
@@
Appendix B SFAL Algorithms
Algorithm 4 SFAL of SSA
Output $\theta^{T}$ given by the server
1:  for $t=0,1...,T-1$ do
2:     $\theta^{t+1}_{i,0}=\theta_{t},\forall i=1,..,m$
3:     for $k=0,1,...,K-1$ do
5:        Accumulate the loss $\mathcal{L}_{loss}^{i.t}$ in the last local epoch
6:     end for
9:  end for
Algorithm 5 SFAL of RSA
Output $\theta^{T}$ given by the server
1:  for $t=0,1...,T-1$ do
2:     $\theta^{t+1}_{i,0}=\theta_{t},\forall i=1,..,m$
3:     for $k=0,1,...,K-1$ do
4:        Sample $u_{i,k_{q}}\in\mathbb{P}^{d}$ uniformly from a unit sphere in $\mathbb{P}^{d}$.
6:        Accumulate the loss $\mathcal{L}_{loss}^{i.t}$ in the last local epoch
7:     end for
10:  end for
Algorithm 6 SFAL of OPSA.
Output $\theta^{T}$ given by the server
1:  for $t=0,1...,T-1$ do
2:     $\theta^{t+1}_{i,0}=\theta_{t},\forall i=1,..,m$
3:     for $k=0,1,...,K-1$ do
5:        Accumulate the loss $\mathcal{L}_{loss}^{i.t}$ in the last local epoch
6:     end for
9:  end for
Appendix C Experiment Setup And Additional Experiments
Figure 4: Generalization Gap of the number of noise $Q$ on SVHN. ($m=40,a=2.0,\rho=2.0$)
Imapct on $Q$. As shown in Figure 4, we can see that adjusting $Q$ significantly reduces the generalization error. This supports our view that more perturbations to the model bring the actual weight space closer to the ideal weight space. In practice, we recommend setting $Q\leq 5$ to reduce computational pressure on each client while still achieving good generalization error.
More Experiments on CIFAR10
Figure 5: Generalization Gap of the attack strength $\rho$ on CIFAR10. ($m=40,a=2.0$)
Figure 6: Generalization Gap of the skew parameter $a$ on CIFAR10. ($m=40,\rho=1.0$)
Figure 7: Generalization Gap of the number of client $m$ on CIFAR10. ($a=1.0,\rho=1.0$)
Appendix D Useful Lemmas
@@
$\|\nabla R_{i}(\theta_{t})-\nabla R(\theta_{t})\|\leq(\xi+6L)D_{i},$
@@
Let $h$ be the adversarial surrogate loss defined in LABEL:adversiarl_surrogate_loss and $\ell$ satisfies Assumption 1. $\forall\theta_{1},\theta_{2}$ and $\forall z\in\mathcal{Z}$, the following hold.
@@
$h(\theta_{1})-h(\theta_{2})\leq\nabla h(\theta^{\prime})^{T}(\theta_{1}-\theta_{2})+\frac{\beta}{2}\|\theta_{1}-\theta_{2}\|^{2}+\xi\|\theta_{1}-\theta_{2}\|.$
$\|G_{\alpha,z}(\theta_{1})-G_{\alpha,z}(\theta_{2})\|\leq(1+\eta\beta)\|\theta-\theta^{\prime}\|+\eta\xi.$
@@
Appendix E Generalization Analyses of VFAL
@@
There exists $\sigma>0$ such that for any $\theta$ and $i\in[m],\mathbb{E}_{z_{i}}[\nabla h(\theta,z_{i})]=\nabla R_{i}(\theta)$ and $\mathbb{E}\|\nabla h(\theta,z)-\nabla R_{i}(\theta)\|^{2}\leq\sigma^{2}$
@@
For simplicity, we use $g(\theta)$ to denote stochastic gradient of loss function $h$.
@@
$=$ $-\frac{K\eta_{t}}{2}\mathbb{E}\|\nabla R(\theta_{t})\|^{2}-\frac{\eta_{t}}{2m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\nabla R_{i}(\theta^{t+1}_{i,k})\|^{2}+\frac{\eta_{t}\beta^{2}}{m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\theta^{t+1}_{i,k}-\theta_{t}\|^{2}+K\xi^{2}\eta_{t},$
@@
$\leq$ $\frac{\beta K\eta_{t}^{2}\sigma^{2}}{m}+\dfrac{\beta K\eta_{t}^{2}}{m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\nabla R_{i}(\theta^{t+1}_{i,k})\|^{2}.$
@@
$+\frac{\beta K\eta_{t}^{2}\sigma^{2}}{m}-\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\dfrac{\eta_{t}}{2m}\left(1-2\beta K\eta_{t}\right)\mathbb{E}\|\nabla R_{i}(\theta_{i,k}^{t+1})\|^{2}.$
@@
$\frac{1}{m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\theta^{t+1}_{i,k}-\theta_{t}\|^{2}\leq\frac{9}{2}K^{2}\eta_{t}^{2}\sigma^{2}+6K^{3}\eta_{t}^{2}\xi^{2}+\sum\limits_{i=1}^{m}\frac{3}{m}K^{3}\eta_{t}^{2}(\xi+6L)D_{i}^{2}+3K^{3}\eta_{t}^{2}\mathbb{E}\|\nabla R(\theta_{t})\|^{2}.$ (12)
@@
$\frac{1}{m}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\mathbb{E}\|\theta^{t+1}_{i,k}-\theta_{t}\|^{2}\leq\frac{9}{2}K^{2}\eta^{2}\sigma^{2}+6K^{3}\eta^{2}\xi^{2}+\sum\limits_{i=1}^{m}\frac{3}{m}K^{3}\eta^{2}(\xi+6L)^{2}D_{i}^{2}+3K^{3}\eta^{2}\mathbb{E}\|\nabla R(\theta_{t})\|^{2}.$
@@
Suppose Assumption 1 and 4 hold, $h(\theta)$ is a $\xi$-approximately $\beta$-gradient Lipschitz and non-convex function w.r.t $\theta$.
$\mathbb{E}\|\theta_{i,k}-\theta_{t}\|\leq(1+\beta\eta_{t})^{K-1}K\eta_{t}(\xi+\sigma+(\xi+6L)D_{i}+\mathbb{E}\|\nabla R(\theta_{t})\|).$
@@
$\leq$ $(1+\beta\eta_{t})\mathbb{E}\|\theta_{i,k}-\theta_{t}\|+\eta_{t}(\xi+\sigma+(\xi+6L)D_{i}+\mathbb{E}\|\nabla R(\theta_{t})\|),$
@@
Suppose Assumption 1 and 4 hold, $h(\theta)$ is a $\xi$-approximately $\beta$-gradient Lipschitz and non-convex function w.r.t $\theta$, for $\eta_{t}\leq\frac{1}{K\beta(t+1)}$ , we have
@@
$+$ $e\sqrt{6\log T\Delta+6\log T(\xi^{2}+\xi L)+\dfrac{\log T\sigma^{2}\pi^{2}}{mK\beta}+\log T\zeta(3)c^{3}\beta^{2}(\frac{27\sigma^{2}}{K\beta}+\frac{36\xi^{2}}{\beta}+\frac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta})}.$
@@
$\leq$ $(1+e)\left[\underbrace{\dfrac{K}{m}\sum\limits_{t=0}^{T-1}\sum\limits_{i=1}^{m}\eta_{t}(\sigma+\xi+(\xi+6L)D_{i})}_{\Lambda_{1}}+\underbrace{\dfrac{K}{m}\sum\limits_{t=0}^{T-1}\sum\limits_{i=1}^{m}\eta_{t}\mathbb{E}\|\nabla R(\theta_{t})\|}_{\Lambda_{2}}\right],$
@@
$\Lambda_{1}\leq$ $\dfrac{K}{m}\sum\limits_{t=0}^{T-1}\sum\limits_{i=1}^{m}\eta_{t}(\sigma+\xi+(\xi+6L)D_{i})\leq\log T(\sigma+\xi+(\xi+6L)D_{\max}).$
@@
$+\dfrac{36\xi^{2}}{\beta(t+1)^{3}}+\dfrac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta(t+1)^{3}}.$
@@
$=$ $\sqrt{6\log T\Delta+6\log T(\xi^{2}+\xi L)+\dfrac{\log^{2}T\sigma^{2}\pi^{2}}{mK\beta}+\log T\zeta(3)c^{3}\beta^{2}(\frac{27\sigma^{2}}{K\beta}+\frac{36\xi^{2}}{\beta}+\frac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta})},$
@@
$+$ $(1+e)\sqrt{6\log T\Delta+6\log T(\xi^{2}+\xi L)+\dfrac{\log T\sigma^{2}\pi^{2}}{mK\beta}+\log T\zeta(3)c^{3}\beta^{2}(\frac{27\sigma^{2}}{K\beta}+\frac{36\xi^{2}}{\beta}+\frac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta})}.$
@@
$\leq$ $(1+\beta\eta_{t})\mathbb{E}\|\theta_{j,k}-\theta_{j,k}^{\prime}\|+\eta_{t}\xi,$
@@
$=$ $e^{\beta K\eta_{t}}(\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+K\eta_{j,t}\xi),$
@@
$\leq$ $(1+\beta\eta_{t})\|\theta_{i,k}-\theta_{i,k}^{\prime}\|+\eta_{t}\xi,$
@@
$\leq$ $(1+\beta\eta_{t})\|\theta_{i,k}-\theta_{i,k}^{\prime}\|+\eta_{t}\xi+\eta_{t}\|g_{i}(\theta_{i,k}^{\prime})-g_{i}^{\prime}(\theta_{i,k}^{\prime})\|,$
@@
$\leq$ $(1+\beta\eta_{t})\mathbb{E}\|\theta_{i,k}-\theta_{i,k}^{\prime}\|+\eta_{t}\xi+\frac{2\eta_{t}}{n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|.$
@@
$=$ $e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+K\eta_{t}e^{\beta K\eta_{t}}(\xi+\dfrac{2}{n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|).$
@@
$\leq$ $\dfrac{1}{m}\sum\limits_{j=1}^{m}e^{\beta K{\eta}_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+\dfrac{1}{m}\sum\limits_{j=1}^{m}e^{\beta K{\eta}_{t}}K\eta_{t}\xi+\dfrac{}{}+\dfrac{2K{\eta}_{t}e^{\beta K\eta_{t}}}{mn_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|.$
@@
$\leq$ $e^{\frac{1}{(1+t)}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+e^{\frac{1}{(1+t)}}\sum\limits_{j=1}^{m}\frac{K\eta_{t}\xi}{m}+e^{\frac{1}{(1+t)}}\sum\limits_{i=1}^{m}\dfrac{2K\eta_{t}}{m^{2}n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|.$
@@
$+$ $\frac{2\beta eT}{mn_{i}}\sqrt{6\log T\Delta+6\log T(\xi^{2}+\xi L)+\dfrac{\log T\sigma^{2}\pi^{2}}{mK\beta}+\log T\zeta(3)c^{3}\beta^{2}(\frac{27\sigma^{2}}{K\beta}+\frac{36\xi^{2}}{\beta}+\frac{18(\xi+6L)^{2}D_{\max}^{2}}{\beta})}.$
@@
$=$ $\mathcal{O}\left(\rho T\log T+\frac{T\sqrt{\log T\Delta}}{mn_{\mathrm{min}}}+\frac{T\log T(\rho+1)D_{\mathrm{max}}}{mn_{\mathrm{min}}}\right).$
@@
(Lin, Zheng, and Jordan (2022)Under assumption 1 and let $\ell_{\gamma}(\theta)=\mathbb{E}_{u\sim\mathbb{P}}[\ell_{\rho}(\theta+\gamma u)]$ where $\mathbb{P}$ is an uniform distribution on an unit ball in $\mathcal{L}_{2}$ norm. We have
$\|\ell_{\gamma}(\theta)-\ell_{\rho}(\theta)\|\leq\gamma L,$
$\|\ell_{\gamma}(\theta)-\ell_{\gamma}(\theta^{\prime})\|\leq L\|\theta-\theta^{\prime}\|,$
$\|\nabla\ell_{\gamma}(\theta)-\nabla\ell_{\gamma}(\theta^{\prime})\|\leq\frac{cL\sqrt{d}}{\gamma}\|\theta-\theta^{\prime}\|,$
@@
$=$ $L\|\theta-\theta^{\prime}\|,$
@@
$\leq$ $\frac{L}{\mathrm{Vol}(\mathbb{B}_{\gamma}(0))}\left(\mathrm{Vol}(\mathbb{B}_{\gamma}(\theta)\setminus\mathbb{B}_{\gamma}(\theta^{\prime}))+\mathrm{Vol}(\mathbb{B}_{\gamma}(\theta^{\prime})\setminus\mathbb{B}_{\gamma}(\theta))\right)$
@@
$=$ $\frac{2L}{c_{d}\gamma^{d}}\mathrm{Vol}(I),$
@@
$\mathrm{Vol}(I)=\mathrm{Vol}(\mathbb{B}_{\gamma}(0))-2V_{cap}(\frac{1}{2}\|\theta-\theta^{\prime}\|)=c_{d}\gamma^{d}-2V_{cap}(\frac{1}{2}\|\theta-\theta^{\prime}\|).$
@@
$\leq$ $(1+\frac{cL\sqrt{d}}{\gamma}\eta)\|x-y\|.$
@@
$\leq$ $(1+\eta)\frac{cL\sqrt{d}}{\gamma}\|\theta_{1}-\theta_{2}\|+2\eta v,$
@@
If the Assumptions 1 and 4 hold and $\eta_{t}\leq\frac{\gamma}{6KcL\sqrt{d}}$, considering $\theta_{i,k+1}=\theta_{i,k}-\eta_{t}\nabla\ell_{\gamma,i}(\theta_{i,k})$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|R(\theta_{t})\|^{2}\leq\mathcal{O}(\frac{\sqrt{\sigma}}{\sqrt{mKT}}+\frac{((\xi+1)D_{\max})^{\frac{2}{3}}}{T^{\frac{2}{3}}}+\frac{\Delta}{T}).$
@@
Suppose the Assumptions 1 and 4 hold, $\ell_{\gamma}(\theta)$ is a $\frac{cL\sqrt{d}}{\gamma}$-gradient Lipschitz and non-convex function w.r.t $\theta$.
@@
Theorem 3. Let $c\geq 0$ be a constant and the step size be chosen as $\eta_{t}\leq\frac{\gamma}{4K\sqrt{d}cL(t+1)}$. Under Assumption 1 and 4, the generalization bound $\varepsilon_{gen}$ with the randomized smoothness approximation satisfies:
@@
$\leq$ $(1+\frac{cL\sqrt{d}}{\gamma}\eta_{t})\mathbb{E}\|\theta_{j,k}-\theta_{j,k}^{\prime}\|+2\eta_{t}\mathbb{E}[v],$
@@
$=$ $e^{\frac{cL\sqrt{d}}{\gamma}K\eta_{t}}(\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+2K\eta_{t}\mathbb{E}[v]),$
@@
$\leq$ $(1+\frac{cL\sqrt{d}}{\gamma}\eta_{t})\|\theta_{i,k}-\theta_{i,k}^{\prime}\|+2\eta_{t}\mathbb{E}[v],$